OpenStudy (anonymous):

Will fan and medal. A bag contains 10 white golf balls and 6 striped golf balls. A golfer wants to add 112 golf balls to the bag. He wants the ratio of white striped golf balls to remain the same. How many of each should he add? Please help! Thanks so much. :)

OpenStudy (anonymous):

The ratio would be 80:48 so he should add 70 white gold balls and 42 striped balls.

OpenStudy (anonymous):

Okay so at the moment the ratio is 10:6 (white:striped) and as we are adding more to the mix, we want to add these with the same ratio. 112/(10+6)=7. So we want 7*10=70 white balls and 7*6=42 striped balls.